OpenStudy (anonymous):

(3,3)(-8,9) into standard form!!!!

OpenStudy (anonymous):

Huh? Do you mean the standard-form equation of the line passing through those points? \[ (x_1,y_1)=(3,3)\\(x_2,y_2)=(-8,9)\\ m=\frac{y_2-y_1}{x_2-x_1}=\frac6{-11}=-\frac6{11}\\ y-y_1=m(x-x_0)\\ y-3=-\frac6{11}(x-3)\\ y=-\frac6{11}x+\frac{18}{11}+\frac{33}{11}\\ y+\frac6{11}x=\frac{51}{11}\\ 11y+6x=51 \]